OpenStudy (anonymous):

Graph each equation. x/3 - y/6 = 1

OpenStudy (anonymous):

\[x/3 - y/6 = 1\] Multiply by 6 \[2x - y = 1\] Move x \[-y = -2x+1\] Multiply by -1 \[y = 2x - 1\] Intercepts = (0,1/2),(-1,0) 

OpenStudy (anonymous):

thank you.